OpenStudy (m_m13):

What is the value of x? will medal 1. x+4=-9 2. 12- x=7 3. -5x=-35 4. x/4-5=6 5. 3(x-2)+5x=18 6. 8x-9=3x=16 7. -2/7+6+1/7x+18

OpenStudy (pawanyadav):

Show your work first.......for all

OpenStudy (anonymous):

I'll do one. Lets say you want to find x for the equation: x+10=5 You can add or subtract the same things on both sides so: x+10-10=5-10 and you can simplify equation so: x=-5

OpenStudy (anonymous):

if you're posting because you need help checking... DO NOT ABUSE THIS: wolframalpha.com that site can check your problems. just input the equation for the roots.

OpenStudy (anonymous):

1) subtract 4 from both side x=-9-4 simplify -9 -4 to -13 x=-13